A houseboat on a river traveled 48km upstream in 6h. It took the houseboat only 4h to make the same trip downstream.
a) How fast would the houseboat have traveled in still water?
b) How fast was the river's current?

Answer by jorel1380(3719)

Let b be the speed of the boat in still water, and c be the speed of the current. Then:
48/b+c=4
48/b-c=6
So:
4b+4c=48
6b-6c=48
12b+12c=144
12b-12c=96
24b=240
b=10
c=2
The boat travels 10 kph in still water, and the current is 2 kph. ☺☺☺☺